OpenStudy (anonymous):

Solve both in substitution and elimination: HELP!!! 3x+2y=8 2x+y=6 Graph you results.

OpenStudy (adilalvi):

multiply 2x+y=6 by 2 and subtruct from equation 1 x will be equal to 2 now sustitute x=2 in equation 1 y=3%2

OpenStudy (anonymous):

That's wrong. 3x+2y=8 2x+y=6 Substitution Method: 2x+y=6 -2x -2x ----------------- y=-2x+6 Plug that in to 3x+2y=8 So, 3x+2(-2x+6)=8 3x-4x+12=8 ---------------- -1x+12=8 -12 -12 --------------- -4/-1x x=4

OpenStudy (anonymous):

The you plug "x" into the top equation: 3(4)+2y=8 12+2y=8 -12 -12 -------------- -4/2y y=-2 Answer: (4,-2)